# ERASER: Machine Unlearning in MLaaS via an Inference Serving-Aware Approach

### v2

@@ Abstract.
In this paper, we propose the ERASER framework for machinE unleaRning in MLaAS via an inferencE seRving-aware approach. ERASER strategically choose appropriate unlearning execution timing to address the inference service obsolescence issue. A novel inference consistency certification mechanism is proposed to avoid the violation of RTBF principle caused by postponed unlearning executions, thereby mitigating the undesirable exposure vulnerability. ERASER offers three groups of design choices to allow for tailor-made variants that best suit the specific environments and preferences of various MLaaS systems. Extensive empirical evaluations across various settings confirm ERASER’s effectiveness, e.g., it can effectively save up to 99% of inference latency and 31% of computation overhead over the inference-oblivion baseline.
copyright: nonecopyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
@@ 2.2. Machine Unlearning
Machine unlearning is the task of removing a specific training sample from the trained ML model. The most basic yet inefficient approach is “retraining-from-scratch”, which retrains the ML model on the remaining training set (i.e., the original training set excluding the sample to be unlearned) from scratch. To avoid the prohibitive computational cost of “retraining-from-scratch”, many efficient machine unlearning mechanisms have been proposed, which can be roughly categorized into exact unlearning mechanisms (Cao and Yang, 2015; Ginart et al., 2019; Brophy and Lowd, 2021; Bourtoule et al., 2021; Ullah et al., 2021; Schelter et al., 2021; Chen et al., 2022c, b; Koch and Soll, 2023) and approximate unlearning mechanisms (Golatkar et al., 2020a; Wu et al., 2020; Golatkar et al., 2020b; Nguyen et al., 2020; Neel et al., 2021; Peste et al., 2021; Golatkar et al., 2021; Warnecke et al., 2021; Izzo et al., 2021; Mahadevan and Mathioudakis, 2021). Exact machine unlearning mechanisms produce models identical to the one produced by “retraining-from-scratch”, therefore capable of completely removing the requested data. Approximate unlearning mechanisms, on the other hand, trade exactness in data removal for improved computational and memory efficiency.
@@ 3.1. Threat Model
Server. In this paper, the server represents a service provider offering MLaaS. The server operates in two phases in existing MLaaS literature. During the training phase, the server builds the ML model using the training dataset that contains personal and potentially sensitive information from data owners. During the inference serving phase, the server processes queriers’ inference requests by feeding the inference sample to the ML models and returning the inference result to the querier. To adhere to the RTBF regulations, the server further takes into account unlearning requests from data owners during the inference serving phase. The server also needs to erase newly identified poisoning data or unauthorized copyrighted data. The server processes the unlearning request with an unlearning mechanism to remove the requested data from the deployed model.
@@ 3.3. Desiderata for Unlearning in MLaaS
There are also other opportunities for improvement in both the security and efficiency of the machine unlearning scheme. However, these areas are orthogonal to the research presented in this paper. Further details can be found in Appendix A.
@@ 4.1. Intuition and Overview
With this selection principle, the challenge is how to certify that the inference result is consistent on models with and without unlearning execution of the pending unlearning requests. Apparently, it is infeasible to compare the inference results by actually executing unlearning, as this would defeat our goal of avoiding its execution in the first place. To certify inference consistency without actual machine unlearning execution, we resort to the concept of model robustness with respect to small changes in the training dataset (Levine and Feizi, 2021; Wang et al., 2022; Jia et al., 2021). Roughly speaking, this refers to the property that two ML models trained on two datasets which differ only on a few data have the same inference results. The ideal robustness entails two favoring consequences: 1) more subsequent inference requests can be responded to before processing the pending unlearning requests, reducing response latency; 2) more unlearning requests can be accumulated and executed all at once by a single unlearning execution, reducing unlearning execution time.
@@ 4.3. Certified Inference Consistency
The proof is relegated to Appendix B. The algorithmic workflow of the inference consistency certification mechanism is depicted in Algorithm 2 at Appendix G.2, while Figure 3 provides an illustration of different cases within the mechanism. Algorithm 1 in Appendix G.1 summarizes the training and inference within ERASER.
@@ 5.1. Design Options
Design Option II: Timing of unlearning. We explore when to execute the pending unlearning requests and provide three options below depending on whether the unlearning requests are processed immediately or waiting until there are inference requests that cannot satisfy the prediction consistency condition.
  • Design Option II-A: ERASER opts to accumulate unlearning requests until an inference request fails to meet the inference consistency condition. At that point, all pending unlearning requests are batch- unlearned. Subsequent inference requests submitted during unlearning can only be responded to if they meets the inference consistency condition based on available model copies. This option is referred to as uncertification-triggered unlearning.
@@ 5.2. Variants of ERASER
Option B-A-B: Single Context for Uncertification-Triggered Unlearning with Postponed Certified Inference (SUTP). SUTP maintains a single context at a time and accumulates unlearning requests until an inference request fails to satisfy the certified inference consistency condition. Instead of updating immediately, constituent models corresponding to unlearning requests are placed on an unlearning waiting list. Once uncertification occurs, the server switches to the unlearning context and executes the unlearning update. Uncertified prediction requests will be postponed until the server completes the unlearning update and switches back to the inference context, as well as the inference requests that arrive during the unlearning context. This variant saves computation overhead compared to DIMP, but at the cost of increased inference latency.
@@
These different variants achieve a tradeoff between privacy, resource consumption, and inference latency. All double-context variants consume more resources to reduce inference latency. DIMP has the lowest inference latency among all variants, but its computation overhead is the highest. The computation overhead of DTTP is the lowest, but its inference latency is the highest. The inference latency and computation overhead of DUTP are between the first two variants. The inference latency of DTTU is close to that of DUTP, but the computation overhead is lower, at the cost of introducing privacy risk. Ultimately, the choice of variant depends on the specific requirements and constraints of a given application.
@@ 5.3. Analysis
Theorem 5.1 shows that $\mathbb{E}(w_{sisa})$ is positively correlated with the frequency of unlearning requests $\frac{n_{u}}{T}$ and the retraining time $r$. Please refer to Appendix C for proofs.
@@
Given the complexity of the actual expression for $\mathbb{E}(w_{dimp})$ (as detailed in Appendix D), which makes direct comparison with $\mathbb{E}(w_{sisa})$ challenging, we present its theoretical upper bound in a form analogous to $\mathbb{E}(w_{sisa})$, which can also be expressed as $\mathbb{E}(w_{dimp})\leq p_{uc}\cdot\mathbb{E}(w_{sisa})$. This expression shows that the processing speed of DIMP is at least $\frac{1}{p_{uc}}$ times that of SISA. Our experimental results indicate that the $p_{uc}$ value is typically less than 0.01, meaning that the waiting time for DIMP does not exceed 1% of SISA. Additional discussion of $p_{uc}$ can be seen in Appendix E.3.
@@ 6.3. Parameters Analysis
In general, as the number of shards increases, inference latency decreases. This is because given the same number of unlearning requests, having more constituent models reduces the proportion of models needing updates, thus increasing the probability of meeting the certified consistency condition. However, when there are too many shards (e.g. more than 30), the decrease in AWT becomes less significant. This is attributed to the fact that each constituent model would have lower accuracy, as shown in Figure 10 at Appendix I.1, resulting in more uncertified prediction results.

### v3

@@ Abstract.
In this paper, we propose the ERASER framework for machinE unleaRning in MLaAS via an inferencE seRving-aware approach. ERASER strategically chooses appropriate unlearning execution timing to address the inference service obsolescence issue. A novel inference consistency certification mechanism is proposed to avoid the violation of RTBF principle caused by postponed unlearning executions, thereby mitigating the undesirable exposure vulnerability. ERASER offers three groups of design choices to allow for tailor-made variants that best suit the specific environments and preferences of various MLaaS systems. Extensive empirical evaluations across various settings confirm ERASER’s effectiveness, e.g., it can effectively save up to 99% of inference latency and 31% of computation overhead over the inference-oblivion baseline.
Machine Unlearning; Machine Learning as a Service
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3670398isbn: 979-8-4007-0636-3/24/10ccs: Security and privacyccs: Computing methodologies Machine learning
@@ 1. Introduction
[…]
The figure shows a server collecting various types of data from data owners. This data is used to train machine learning models, represented by a neural network diagram. Companies like Google, OpenAI, and Microsoft utilize these models to process inference requests from queriers and provide responses. There is also a process for handling unlearning requests to update the models as needed for privacy and other concerns.
@@
[…]
This figure compares the ERASER strategy with naive strategies in the context of machine learning models and their handling of inference and unlearning requests. On the left side, the training set consists of various images. This training set is split into multiple subsets, each used to train separate neural network models. The process involves training the models with these subsets and then using them for predictions. The right side of the figure illustrates the handling of inference and unlearning requests. The comparison is divided into three parts: 1. SISA-Unlearning Request First: This section shows how inference and unlearning requests are handled by the SISA strategy, prioritizing unlearning requests (UR) before inference requests (IR). Green bars indicate inference without privacy risk, red bars indicate inference with privacy risk, and yellow bars signify unlearning execution. 2. SISA-Inference Request First: This section depicts the handling of requests when inference requests are prioritized before unlearning requests. Similar to the previous section, green, red, and yellow bars indicate inference without privacy risk, inference with privacy risk, and unlearning execution, respectively.3. Eraser: This final section represents a naive strategy where inference requests are processed without considering privacy risks or unlearning executions.The comparison highlights the effectiveness of the SISA strategy in managing privacy risks associated with inference requests, as indicated by the varying lengths and colors of the bars in each section.
@@
Due to page limits, please refer to https://arxiv.org/abs/2311.16136 for further discussions, detailed proofs and additional experiments.
@@ 2.2. Machine Unlearning
Machine unlearning is the task of removing a specific training sample from the trained ML model. The most basic yet inefficient approach is “retraining-from-scratch”, which retrains the ML model on the remaining training set (i.e., the original training set excluding the sample to be unlearned) from scratch. To avoid the prohibitive computational cost of “retraining-from-scratch”, many efficient machine unlearning mechanisms have been proposed, which can be roughly categorized into exact unlearning mechanisms (Cao and Yang, 2015; Ginart et al., 2019; Brophy and Lowd, 2021; Bourtoule et al., 2021; Ullah et al., 2021; Schelter et al., 2021; Chen et al., 2022c, b; Koch and Soll, 2023; Wang et al., 2023; Tao et al., 2024) and approximate unlearning mechanisms (Golatkar et al., 2020a; Wu et al., 2020; Golatkar et al., 2020b; Nguyen et al., 2020; Neel et al., 2021; Peste et al., 2021; Golatkar et al., 2021; Warnecke et al., 2021; Izzo et al., 2021; Mahadevan and Mathioudakis, 2021; Liu et al., 2023, 2024; Lin et al., 2023). Exact machine unlearning mechanisms produce models identical to the one produced by “retraining-from-scratch”, therefore capable of completely removing the requested data. Approximate unlearning mechanisms, on the other hand, trade exactness in data removal for improved computational and memory efficiency.
@@ 3.1. Threat Model
Server. In this paper, the server represents a service provider offering MLaaS. The server operates in two phases in existing MLaaS literature. During the training phase, the server builds the model using the training dataset that contains personal and potentially sensitive information from data owners. During the inference serving phase, the server processes queriers’ inference requests by feeding the inference sample to the models and returning the result to the querier. To adhere to the RTBF regulations, the server further takes into account unlearning requests from data owners during the inference serving phase. The server also needs to erase newly identified poisoning data or unauthorized copyrighted data.
@@ 3.3. Desiderata for Unlearning in MLaaS
What we do not achieve in this paper. There are also other opportunities for improvement in both the security and efficiency of the machine unlearning scheme. However, these areas are orthogonal to the research presented in this paper.
Regarding extra privacy risk: We focus on ensuring that data owners who have submitted unlearning requests do not suffer from extra privacy risk due to the postponed unlearning execution. However, we do not address other sources of privacy threats, such as general privacy attacks considered in MIA literature where any data owners can be the victim (Shokri et al., 2017; Salem et al., 2019; Li and Zhang, 2021; Liu et al., 2022a).
@@ 4.1. Intuition and Overview
With this selection principle, the challenge is how to certify that the inference result is consistent on models with and without unlearning execution of the pending unlearning requests. Apparently, it is infeasible to compare the inference results by actually executing unlearning, as this would defeat our goal of avoiding its execution in the first place. To certify inference consistency without actual machine unlearning execution, we resort to the concept of model robustness with respect to small changes in the training dataset (Levine and Feizi, 2021; Wang et al., 2022; Jia et al., 2021). Roughly speaking, this refers to the property that two ML models trained on two datasets which differ only on a few data have the same inference results. This naturally evokes the idea of utilizing Differential Privacy (DP) (Dwork et al., 2006) techniques for model training. However, the primary challenge with the DP-based approach is that the robustness offered by DP allows only for approximate unlearning, rather than exact unlearning. As the number of unlearning request keeps increasing during the MLaaS service, the privacy bound of DP may be breached with the growth of to-be-erased data. Additionally, the noise introduced by DP can significantly decrease the model’s utility. The ideal robustness entails two favoring consequences: 1) more subsequent inference requests can be responded to before processing the pending unlearning requests, reducing response latency; 2) more unlearning requests can be accumulated and executed all at once by a single unlearning execution, reducing unlearning execution time.
@@
1:Training dataset $\bm{\mathscr{D}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$
2:
5:Launch inference service with $F=\{f_{1},\dots,f_{K}\}$
6:
7:for $t\in[1,2,\dots]$ do
9:         Record it as a pending unlearning request
13:         if satisfy certified inference consistency condition  then
15:         else
16:             Process in accordance with specific design options.
17:         end if
18:    end if
19:end for
20:Responses for inference and unlearning requests.
@@
[…]
This figure illustrates three representative cases for the certified inference consistency mechanism in a machine learning model. The left part of the figure shows an inference request (IR) for an image of a husky, which is processed by multiple models, each represented by neural network icons. Some models are marked with red question marks, indicating uncertainty in their predictions. Case 1: The prediction result is Dog: 5, Cat: 3, Panda: 1. A potential flip is considered, and for the Cat and Panda classes, the inequalities 2y_1 + y_3 ¡ Γare checked with different values of Γ. The worst condition results in Dog: 5, Cat: 4, Panda: 0. The case is determined to be certified. Case 2: The prediction result is Dog: 5, Cat: 3, Panda: 1. A potential flip is considered, and the inequality 2y_1 + y_3 ¿ Γis checked for the Cat class, resulting in Dog: 4, Cat: 5, Panda: 0 under the worst condition. This case is determined to be uncertified. Case 3: The prediction result is Dog: 4, Cat: 3, Panda: 2. A potential flip is considered, and the inequalities for the Cat and Panda classes are checked, resulting in Dog: 4, Cat: 0, Panda: 5 under the worst condition. This case is also determined to be uncertified. Each case involves checking conditions to determine if the prediction remains consistent under potential flips, with results categorized as certified or uncertified based on the consistency check outcomes.
@@ 4.3. Certified Inference Consistency
Algorithm 2 Inference Consistency Certification Mechanism
2:$certification_{\tt{q}}=True$
3:for $k\in[1,2,\dots,K]$ do
4:    Inference with constituent model: $f_{k}^{\tt O}(\mathbf{z})$
5:    if $\bm{\mathscr{S}}_{k}^{\tt O}\cap\tt{U}_{k}^{t}\neq\varnothing$  then
6:         Mark as $\bm{\mathscr{S}}_{k}^{t}\neq\bm{\mathscr{S}}_{k}^{\tt O}$
7:    end if
8:end for
10:for $\forall y_{b}\in\mathbb{Y},y_{b}\neq y_{a}$ do
15:    if $2\gamma_{1}+\gamma_{3}>\Gamma_{b}$ then
16:         $certification_{\tt{q}}=False$
17:         Break
18:    end if
19:end for
20:if $certification_{\tt{q}}$ then
22:else
23:    The consistency of $\tt{q}$ can not be certified
24:end if
25:$certification_{\tt{q}}$
All the proofs in this paper is relegated to Appendices in the arXiv version. Figure 3 provides an illustration of different cases within the mechanism. Algorithm 2 summarises the inference consistency certification mechanism.
4.4. Extension to Different Unlearning Methods
Although the design of ERASER  draws inspiration from SISA, which retrains the constituent models from scratch, ERASER  can also be extended beyond SISA. Any exact-unlearning algorithm applicable to the target model can be utilized within the ERASER scheme as a substitute for “retraining from scratch” to update constituent models. For $(\epsilon,\delta)$-approximate-unlearning algorithms, the accumulation of multiple unlearning requests might continuously amplify $\epsilon$ and $\delta$. However, we believe that leveraging composition techniques from differential privacy could help address this issue, which will be further explored in our future work.
@@ 5.1. Design Options
Design Option II: Timing of unlearning. We explore when to execute the pending unlearning requests and provide three options below depending on whether the unlearning requests are processed immediately or wait until there are inference requests that cannot satisfy the prediction consistency condition.
  • Design Option II-A: ERASER opts to accumulate unlearning requests until an inference request fails to meet the inference consistency condition. At that point, all pending unlearning requests are batch-unlearned. Subsequent inference requests submitted during unlearning can only be responded to if they meet the inference consistency condition based on available model copies. This option is referred to as uncertification-triggered unlearning.
@@ 5.2. Variants of ERASER
Option B-A-B: Single Context for Uncertification-Triggered Unlearning with Postponed Certified Inference (SUTP). SUTP maintains a single context at a time and accumulates unlearning requests until an inference request fails to satisfy the certified inference consistency condition. Instead of updating immediately, constituent models corresponding to unlearning requests are placed on an unlearning waiting list. Once uncertification occurs, the server switches to the unlearning context and executes the unlearning update. Uncertified prediction requests will be postponed until the server completes the unlearning update and switches back to the inference context, as well as the inference requests that arrive during the unlearning. This variant saves computation overhead compared to DIMP, but at the cost of increased inference latency.
@@
[…]
This figure illustrates four variants of the ERASER framework with double contexts for handling inference requests (IR) and unlearning requests (UR) over a timeline from T1 to T9. DIMP : At T1, a UR arrives and triggers the unlearning process. IRs continue to arrive and are processed with potential delays due to the unlearning process. Unlearning completes by T5. DUTP: At T1, a UR arrives, but unlearning is delayed until T5. Inference requests are processed without unlearning until T5. Unlearning begins at T5 and finishes by T7. DTTU: Inference requests are processed, and unlearning is triggered only when the number of uncertified IRs exceeds a threshold (set to 3) at T6. Unlearning starts at T6 and completes by T8. No re-processing of previous IRs occurs. DTTP: Similar to DTTU, but once unlearning finishes at T8, there is a re-processing phase for previous IRs that were processed before unlearning started. This re-processing occurs at T9. The timeline on the left shows the decision points for unlearning based on the arrival of URs and the certification status of IRs. The process flow within each variant depicts how IRs and URs are managed to maintain model consistency and privacy.
These different variants achieve a tradeoff between privacy, resource consumption, and inference latency. All double-context variants consume more resources to reduce inference latency. DIMP has the lowest inference latency among all variants, but its computation overhead is the highest. The computation overhead of DTTP is the lowest, but its inference latency is the highest. The inference latency and computation overhead of DUTP are between the first two variants. The inference latency of DTTU is close to that of DUTP, but the computation overhead is lower, at the cost of introducing privacy risk. Ultimately, the choice of variant depends on the specific requirements and constraints of a given application. A server can also dynamically adjust among different variants based on the current computational resources and response time to pursue higher resource utilization.
@@ 5.3. Analysis
Theorem 5.1 shows that $\mathbb{E}(w_{sisa})$ is positively correlated with the frequency of unlearning requests $\frac{n_{u}}{T}$ and the retraining time $r$.
@@
Given the complexity of the actual expression for $\mathbb{E}(w_{dimp})$, which makes direct comparison with $\mathbb{E}(w_{sisa})$ challenging, we present its theoretical upper bound in a form analogous to $\mathbb{E}(w_{sisa})$, which can also be expressed as $\mathbb{E}(w_{dimp})\leq p_{uc}\cdot\mathbb{E}(w_{sisa})$.
This expression shows that the processing speed of DIMP is at least $\frac{1}{p_{uc}}$ times that of SISA. Our experimental results indicate that the $p_{uc}$ value is typically less than 0.01, meaning that the waiting time for DIMP does not exceed 1% of SISA. We can also qualitatively analyze $p_{uc}$ based on the characteristics of the model:
@@ 6.1. Experiment Setup
[…]
This figure presents the cumulative percentage of waiting time for different strategies across four datasets: Purchase, SVHN, MNIST, and Imagenette. Each plot shows the cumulative percentage of waiting time (on a logarithmic scale) on the x-axis, against the percentage of requests processed on the y-axis. The strategies compared include SISA, DIMP, SUTP, STTU, DUTP, DTTU, STTP, and DTTP, each represented by different colored lines and markers. Purchase Dataset: The cumulative waiting time for requests shows a significant variation among the strategies, with SISA (orange dashed line) generally exhibiting longer waiting times compared to other strategies, which converge faster. SVHN Dataset: The patterns are similar, where SISA again has the longest waiting times, while other strategies like DUTP (blue dashed line) and DTTU (blue line) show quicker convergence. MNIST Dataset: The waiting times are more uniform across different strategies, with most lines clustering towards the top, indicating faster processing for most requests. Imagenette Dataset: There is a noticeable spread in waiting times, with SISA showing significantly longer times. Strategies like DUTP and DTTU again show better performance in terms of reduced waiting times. The figure highlights the effectiveness of different strategies in managing and reducing the waiting time for processing requests in machine learning systems, with SISA generally showing poorer performance compared to other methods.
@@ 6.3. Parameters Analysis
[…]
This figure evaluates the impact of the number of shards on the average waiting time (AWT) and the number of retraining (NoR) for two datasets, Purchase and SVHN. The figure consists of four plots: Top Left (Impact on AWT for Purchase): This plot shows the average waiting time for different strategies as the number of shards increases from 5 to 50. The strategies include SISA, SUTP, STTU, STTP, DIMP, DUTP, DTTU, and DTTP. SISA generally shows higher waiting times, while strategies like DIMP and DUTP demonstrate significantly lower waiting times as the number of shards increases. Top Right (Impact on NoR for Purchase): This plot displays the number of retraining events required for different strategies as the number of shards increases. SISA consistently shows the highest number of retraining events, while other strategies like DTTP and DTTU show relatively fewer retraining events. Bottom Left (Impact on AWT for SVHN): This plot illustrates the average waiting time for different strategies with the SVHN dataset as the number of shards increases. Similar to the Purchase dataset, SISA has higher waiting times, while DIMP and DUTP show much lower waiting times. Bottom Right (Impact on NoR for SVHN): This plot shows the number of retraining events for different strategies with the SVHN dataset. SISA remains the highest, whereas DTTP and DTTU demonstrate fewer retraining events. Overall, the figure highlights the efficiency of different strategies in terms of reducing average waiting time and the number of retraining events, with DIMP and DUTP performing well compared to SISA.
@@
In general, as the number of shards increases, inference latency decreases. This is because given the same number of unlearning requests, having more constituent models reduces the proportion of models needing updates, thus increasing the probability of meeting the certified consistency condition. However, when there are too many shards (e.g. more than 30), the decrease in AWT becomes less significant. This is attributed to the fact that each constituent model would have lower accuracy, resulting in more uncertified prediction results.
@@
[…]
This figure evaluates the impact of the ratio of unlearning requests on the average waiting time (AWT) and the number of retraining (NoR) for two datasets, Purchase and SVHN. The figure consists of four plots: Top Left (Impact on AWT for Purchase): This plot shows the average waiting time for different strategies as the ratio of unlearning requests increases from 0.01 to 0.5. The strategies include SISA, SUTP, STTU, STTP, DIMP, DUTP, DTTU, and DTTP. SISA generally shows higher waiting times, while strategies like DIMP and DUTP demonstrate significantly lower waiting times as the ratio of unlearning requests increases. Top Right (Impact on NoR for Purchase): This plot displays the number of retraining events required for different strategies as the ratio of unlearning requests increases. SISA consistently shows the highest number of retraining events, while other strategies like DTTP and DTTU show relatively fewer retraining events. Bottom Left (Impact on AWT for SVHN): This plot illustrates the average waiting time for different strategies with the SVHN dataset as the ratio of unlearning requests increases. Similar to the Purchase dataset, SISA has higher waiting times, while DIMP and DUTP show much lower waiting times. Bottom Right (Impact on NoR for SVHN): This plot shows the number of retraining events for different strategies with the SVHN dataset. SISA remains the highest, whereas DTTP and DTTU demonstrate fewer retraining events. Overall, the figure highlights the efficiency of different strategies in terms of reducing average waiting time and the number of retraining events, with DIMP and DUTP performing well compared to SISA.
@@
[…]
This figure evaluates the impact of request density on the average waiting time (AWT) and the number of retraining (NoR) for two datasets, Purchase and SVHN. The figure consists of four plots: Top Left (Impact on AWT for Purchase): This plot shows the average waiting time for different strategies as the request density increases from 0.1 to 10. The strategies include SISA, SUTP, STTU, STTP, DIMP, DUTP, DTTU, and DTTP. SISA generally shows higher waiting times, while strategies like DIMP and DUTP demonstrate significantly lower waiting times as the request density increases. Top Right (Impact on NoR for Purchase): This plot displays the number of retraining events required for different strategies as the request density increases. SISA consistently shows the highest number of retraining events, while other strategies like DTTP and DTTU show relatively fewer retraining events. Bottom Left (Impact on AWT for SVHN): This plot illustrates the average waiting time for different strategies with the SVHN dataset as the request density increases. Similar to the Purchase dataset, SISA has higher waiting times, while DIMP and DUTP show much lower waiting times. Bottom Right (Impact on NoR for SVHN): This plot shows the number of retraining events for different strategies with the SVHN dataset. SISA remains the highest, whereas DTTP and DTTU demonstrate fewer retraining events. Overall, the figure highlights the efficiency of different strategies in terms of reducing average waiting time and the number of retraining events, with DIMP and DUTP performing well compared to SISA.
@@
[…]
This figure evaluates the impact of the threshold of uncertification ratio on the average waiting time (AWT) and the number of retraining (NoR) for two datasets, Purchase and SVHN. The figure consists of two plots: Left (Impact on Purchase): This plot shows the average waiting time and the number of retraining events for different strategies as the threshold of uncertification ratio increases from 0.005 to 0.1. The strategies include STTU, DTTU, STTP, and DTTP. The average waiting time is indicated by lines with circular markers, while the number of retraining events is indicated by lines with square markers. STTU generally shows a higher number of retraining events and a decreasing trend in waiting time, while DTTP shows the lowest number of retraining events with a relatively stable waiting time. Right (Impact on SVHN): This plot illustrates the same metrics for the SVHN dataset. Similar patterns are observed, with STTU showing a higher number of retraining events and decreasing waiting time, whereas DTTP maintains a low number of retraining events and stable waiting times. Overall, the figure highlights how varying the threshold of uncertification ratio affects the performance of different strategies in terms of average waiting time and the number of retraining events, with DTTP performing consistently well compared to others.